OpenStudy (anonymous):

solve triangle ABC given that A=56 degrees, B=57 degrees and b=9 a. C=64 degrees, a=10.2, c=9.8 b. C=67 degrees, a= 10.2, c=9.88 c. C=67 degrees, a=8.9, c=9.88 d. C=66 degrees, a=8.7, c=9.88

OpenStudy (campbell_st):

well you can find C by angle sum of a triangle... then find the sides using the law of sines \[\frac{b}{\sin(B)} = \frac{a}{\sin(A)} = \frac{c}{\sin(C)}\] hope this helps.

OpenStudy (221emily):